OpenStudy (anonymous):

The bottom of a rectangular swimming pool is 13 m × 24 m. If the atmospheric pressure above the swimming pool changes from 744 to 768 mm of mercury, determine the amount by which the force on the bottom of the pool increases? (Assume the density of mercury is 13.6 103 kg/m3.)

OpenStudy (anonymous):

I know P=F/A. Obviously area=312. But i don't understand what to do with this. :(

OpenStudy (anonymous):

i don't know whether i am right or not. but i can suggest you something... so, should i?

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

assuming that the swimming pool is filled with water, the pressure on the bottom will be :- pressure on upper side of pool + pressure due to the water in the pool. which can be calculated by the formula (density * acceleration due to gravity * height)

OpenStudy (anonymous):

so...P1 at top of pool = Patm. P2 at bottom of pool = P1 + h (rho) g?

OpenStudy (anonymous):

yep..!

OpenStudy (anonymous):

How do I convert atmospheric pressure 1.013x10^5N/m^2 to mercurial pressure?

OpenStudy (anonymous):

744mm (Hg)

OpenStudy (anonymous):

its equal to 760 mm of mercury i think...

OpenStudy (anonymous):

1 atm = 760mm oh Hg

OpenStudy (anonymous):

Oh. Thanks. I didn't know I should just "know" that!

OpenStudy (anonymous):

now, can you solve the ques?

OpenStudy (anonymous):

I'll try. THanks for your help.

OpenStudy (anonymous):

welcome!